# Real $K$-Theory for $C^{*}$-Algebras: just the facts October 22, 2025

### v2

@@ 25. $K$-theory and Unitaries
The real $K$-theory groups $KO_{i}(A)$ of a real $C^{*}$-algebra $A$ were defined in Definitions 6.10 and 6.11 in terms of homotopy groups, or in terms of suspensions. We have also seen that it can be defined in terms of Clifford algebras as in Section 18. The picture of $K_{i}(A_{\mathbb{C}})$ and $KO_{i}(A)$ here is equivalent to those definitions but is a distinct picture of $K$-theory. It has the advantage of giving us a way to represent all elements of $KO_{i}(A)$ in terms of a specific (homotopy class of a) unitary in a matrix algebra over $A_{\mathbb{C}}$. This concrete representation allows physicists to develop formulas for topological markers.
@@
Table 1 is reproduced from [18], and it summarizes the unitary picture of $K$-theory that will be described in this section. Each line gives the way of viewing $K_{i}(A_{\mathbb{C}})$ or $KO_{i}(A)$ in terms of unitaries. For example, Propositions 25.1 and 25.2 below are summarized in the lines of Table 1 corresponding to $K_{0}(A_{\mathbb{C}})$ and $KO_{0}(A)$, respectively. The other lines in Table 1 similarly show how to represent $K_{i}(A_{\mathbb{C}})$ and $KO_{i}(A)$ for $i\neq 0$ in this unitary picture, and they correspond to the propositions in this section which state the result explicitly.
Table 1. Unitary Picture of $K$-theory
@@
real $KO_{-1}(A)$ 1 $u^{\tau}=u$ 11
$KO_{0}(A)$ 2 $u=u^{*}$, $u^{\tau}=u^{*}$ $\left(\begin{smallmatrix}{1}&{0}\\ {0}&{-1}\end{smallmatrix}\right)$
@@
Explanation of Table 1. For a real $C^{*}$-algebra $A_{\mathbb{C}}$, define
@@
where $\vee$ denotes disjoint unions. We will discuss the identification of matrices of different sizes later. The way that we do this is not the same for all $i$.
For each row of the table, there is a given value of $\ell$ and the corresponding group $K_{i}(A_{\mathbb{C}})$ or $KO_{i}(A)$ is isomorphic to a group [^85] In each of the ten cases we are taking equivalence classes of certain spaces of matrices. We must show in each case that this has indeed the structure of a group. consisting of equivalence classes of elements in $\mathcal{U}_{\ell,\infty}(A_{\mathbb{C}})$ that satisfy the relation $\mathcal{S}$. More precisely, if $a\in\mathcal{U}_{n}(A_{\mathbb{C}})$ (where $n$ is a multiple of $\ell$) and $a$ satisfies the relation $\mathcal{S}$, then $[a]$ represents an element of $K_{i}(A_{\mathbb{C}})$ or $KO_{i}(A)$. If two elements $a$ and $b$ in $\mathcal{U}_{n}(A_{\mathbb{C}})$ that satisfy $\mathcal{S}$ can be connected by a path $a_{t}$ in $\mathcal{U}_{n}(A_{\mathbb{C}})$ with $a_{0}=a$, $a_{1}=b$ and $a_{t}$ satisfying $S$ for each $t$, then we have $[a]=[b]$ in $K_{i}(A_{\mathbb{C}})$ or $KO_{i}(A)$ (depending on which row of the table we are working in). In each case the group operation for $K_{i}(A_{\mathbb{C}})$ or $KO_{i}(A)$ is given by
@@
Finally the last entry of each row of Table 1 shows the unitary $I\in\mathcal{U}_{\ell,\infty}(A_{\mathbb{C}})$ that represents the trivial group element in $K_{i}(A_{\mathbb{C}})$ or $KO_{i}(A)$. In particular this trivial element gives us the formulas
@@
In what remains of this section, we explain this picture more fully and we state the propositions that form the basis of this picture, and we present some examples. For proofs of these propositions, including how one passes from disjoint unions to unions, see Boersema and Loring [20] Section 5 for the even case and Section 6 for the odd case.
@@
Remarkably, by making a small sign change to the relation $\mathcal{S}$ characterizing $KO_{0}(A)$, we obtain the unitary picture of $KO_{2}(A)$, as in the next Proposition and as shown in the line of Table 1 for $KO_{2}(A)$. We will not give proofs for this proposition as indicated before.
@@
on $M_{2}(A)$, and let $\sharp\otimes\tau$ be the corresponding involution on $M_{2n}(A)=M_{2}({\mathbb{R}})\otimes M_{n}(A)$. Then we have the following propositions. For these propositions, the unitary $I$ such that $[I]$ represents the identity element of $KO_{i}(A)$ is listed in Table 1.
@@
This completes the discussion of the even cases of Table 1. In the odd cases, we first discuss the well-known exponential map in the complex case. The following is Theorem 10.1.3 [69].
@@
We have $K_{1}(A_{\mathbb{C}})=K_{0}(SA_{\mathbb{C}})$ (as presented in Theorem 8.5) and for the moment we let $K_{1}^{u}(A_{\mathbb{C}})$ denote the unitary group referred to in this proposition. It consists of homotopy classes of elements in $\mathcal{U}_{\infty}(A)_{\mathbb{C}})$, subject to the equivalence relation in which a unitary $u\in M_{n}(A_{\mathbb{C}})$ is identified with $\left(\begin{smallmatrix}u&0\\ 0&1\end{smallmatrix}\right)$ in $M_{n+1}(A_{\mathbb{C}})$. The identity is the class of the unit $[1]$ and the group inverse of an element $[u]$ in $K_{1}^{u}(A_{\mathbb{C}})$ is given by $u^{*}]$, since the unitary $\left(\begin{smallmatrix}u&0\\ 0&u^{*}\end{smallmatrix}\right)$ can be connected by a path of unitaries to $\left(\begin{smallmatrix}1&0\\ 0&1\end{smallmatrix}\right)$.
@@
Let $u$ be a unitary in $M_{n}(A_{\mathbb{C}})$. Let $x_{t}$ be a path of unitaries in $\mathcal{U}_{2n}({\mathbb{R}})$ from $\left(\begin{smallmatrix}0&1_{n}\\ 1_{n}&0\end{smallmatrix}\right)$ to $1_{2n}$, and let $v_{t}$ be the path
@@
Then $p_{t}=v_{t}\left(\begin{smallmatrix}1_{n}&0\\ 0&1_{n}\end{smallmatrix}\right)v_{t}^{*}$ is a projection in $C([0,1],M_{2n}(A_{\mathbb{C}}))$ with $p_{0}=\left(\begin{smallmatrix}1_{n}&0\\ 0&0\end{smallmatrix}\right)$ and $p_{1}=\left(\begin{smallmatrix}1_{n}&0\\ 0&0\end{smallmatrix}\right)$. The element
@@
We refer to the proof of Theorem 10.1.3 for the details that $\Theta$ is well-defined and gives an isomorphism, proving the Proposition. ∎
@@
Let $A$ be a real unital $C^{*}$-algebra. Then $KO_{1}(A)$ is isomorphic to the group of homotopy classes in $\mathcal{U}_{\infty}(A_{\mathbb{C}})$ of unitaries $u$ that satisfy $u^{*}=u^{\tau}$.
Note here that the condition $u^{*}=u^{\tau}$ is equivalent to just requiring that $u$ is a unitary in $A\subset A_{\mathbb{C}}$.
@@
This proof is the same as in the complex case, except we note that $\left(\begin{smallmatrix}0&1_{n}\\ 1_{n}&0\end{smallmatrix}\right)$ is not in the same homotopy class of $1_{2n}$ in $M_{2n}({\mathbb{R}})$ when $n$ is odd. However, by moving to a larger matrix algebra they become equivalent.
If $n$ is odd, then
@@
is in the same homotopy class of unitaries as $1_{2n+1}$. Using this, we can produce a path of from $1_{2n+1}$ to $\mathrm{diag}(u,u^{*},1)$ in $M_{2n+1}({\mathbb{R}})$. The rest of the proof is the same as for Proposition 25.6. ∎
@@
Since $S{\mathbb{R}}=C_{o}^{{\mathbb{R}}}({\mathbb{R}})$ is not unital, we must consider the unitization
@@
where $\pi$ is evaluation at $z=1$. The unitaries in the fourth column of the table shows the generators of the various non-trivial $KO_{i}$ groups in this picture.

### v3

@@ 25. $K$-theory, KOKO-theory, and Unitaries
The real $K$-theory groups $KO_{i}(A)$ of a real $C^{*}$-algebra $A$ were defined in Definitions 6.10 and 6.11 in terms of homotopy groups, or in terms of suspensions. We have also seen that it can be expressed in terms of Clifford algebras as in Section 18. The picture of $K_{i}(A_{\mathbb{C}})$ and $KO_{i}(A)$ here is equivalent to those definitions but is a distinct picture of $K$-theory. It has the advantage of giving us a way to represent all elements of $KO_{i}(A)$ in terms of a specific (homotopy class of a) unitary in a matrix algebra over $A_{\mathbb{C}}$. This concrete representation allows physicists to develop formulas for topological markers.
@@
Table 25.1 is reproduced from [18], and it summarizes the unitary picture of $K$-theory that will be described in this section. Each line gives the way of viewing $K_{i}(A_{\mathbb{C}})$ or $KO_{i}(A)$ in terms of unitaries. We will describe here in general terms the information described in the table, but the content of each row is stated explicitly in the Propositions 25.1 through 25.10 which follow. For example, Propositions 25.2 and 25.3 below are summarized in the lines of Table 25.1 corresponding to $K_{0}(A_{\mathbb{C}})$ and $KO_{0}(A)$, respectively. The other lines in Table 25.1 similarly show how to represent $K_{i}(A_{\mathbb{C}})$ and $KO_{i}(A)$ for $i\neq 0$ in this unitary picture.
Table 25.1.
Unitary Picture of $K$-theory: Summary of Propositions 25.1 through 25.10
@@
real $KO_{0}(A)$ 2 $u=u^{*}$, $u^{\tau}=u^{*}$ $\left(\begin{smallmatrix}{1}&{0}\\ {0}&{-1}\end{smallmatrix}\right)$
@@
$KO_{7}(A)$ 1 $u^{\tau}=u$ 11
We establish some notation for sets of matrices with arbitrary size. Let $\vee$ denote disjoint union. For a real $C^{*}$-algebra $A$, define
@@
We will later discuss the identification of matrices of different sizes, but note that the way that we do this is not the same for all $KO_{i}(A)$.
Columns in a nutshell: The value of $\ell$ is $1,2$ or $4$. It indicates whether we are using $M_{1,\infty},M_{2,\infty},$ or $M_{4,\infty}$ size matrices in that row. The entry in Column $\mathcal{S}$ gives us the critical relation that we need on the unitaries that we are using in this row. The entry in column $I$ tells us how the identity in that group is represented. Now here’s some detail.
For each row of the table, there is a given value of $\ell$ and a relation $\mathcal{S}$. The corresponding group $K_{i}(A_{\mathbb{C}})$ or $KO_{i}(A)$ is isomorphic to a group [^88] In each of the ten cases we are taking equivalence classes of certain spaces of matrices. It is necessary to show in each case that this has indeed the structure of a group. consisting of equivalence classes of elements in $\mathcal{U}_{\ell,\infty}(A_{\mathbb{C}})$ that satisfy the relation (or relations) given in the column labeled $\mathcal{S}$.
More precisely, if $a\in\mathcal{U}_{n}(A_{\mathbb{C}})$ (where $n$ is a multiple of $\ell$) and $a$ satisfies the relation $\mathcal{S}$, then $[a]$ represents an element of $K_{i}(A_{\mathbb{C}})$ or $KO_{i}(A)$. For example, the table indicates that $K_{0}(A_{\mathbb{C}})$ is represented by self-adjoint unitaries in $M_{2n}(A_{\mathbb{C}})$ and $K_{1}(A_{\mathbb{C}})$ is represented by arbitrary unitaries in $M_{n}(A_{\mathbb{C}})$, as stated formally in Propositions 25.1 and 25.6 below.
If two elements $a$ and $b$ in $\mathcal{U}_{n}(A_{\mathbb{C}})$ that satisfy $\mathcal{S}$ can be connected by a path $a_{t}$ in $\mathcal{U}_{n}(A_{\mathbb{C}})$ with $a_{0}=a$, $a_{1}=b$ and $a_{t}$ satisfying $S$ for each $t$, then we have $[a]=[b]$ in $K_{i}(A_{\mathbb{C}})$ or $KO_{i}(A)$ (depending on which row of the table we are working in). In each case the group operation for $K_{i}(A_{\mathbb{C}})$ or $KO_{i}(A)$ is given by
@@
Finally the last entry of each row of Table 25.1 shows the unitary $I\in\mathcal{U}_{\ell,\infty}(A_{\mathbb{C}})$ that represents the trivial group element in $K_{i}(A_{\mathbb{C}})$ or $KO_{i}(A)$. In particular this trivial element gives us the formulas
@@
In what remains of this section, we explain this picture more fully and we state the propositions that form the basis of this picture, and we present some examples. For full proofs of these propositions, including how one passes from disjoint unions to unions, see Boersema and Loring [20] Section 5 for the even case and Section 6 for the odd case.
@@
Remarkably, by making a small sign change to the relation $\mathcal{S}$ characterizing $KO_{0}(A)$, we obtain the unitary picture of $KO_{2}(A)$, as in the next Proposition and as shown in the line of Table 25.1 for $KO_{2}(A)$. We will not give proofs for this proposition as indicated before.
@@
on $M_{2}(A)$, and let $\sharp\otimes\tau$ be the corresponding involution on $M_{2n}(A)=M_{2}({\mathbb{R}})\otimes M_{n}(A)$. Then we have the following propositions. For these propositions, the unitary $I$ such that $[I]$ represents the identity element of $KO_{i}(A)$ is listed in Table 25.1.
@@
This completes the discussion of the even cases of Table 25.1. In the odd cases, we first discuss the well-known exponential map in the complex case. The following is Theorem 10.1.3 [69].
@@
We have $K_{1}(A_{\mathbb{C}})=K_{0}(SA_{\mathbb{C}})$ (as presented in Theorem 8.5) and for the moment we let $K_{1}^{u}(A_{\mathbb{C}})$ denote that group of homotopy classes in
$\mathcal{U}_{\infty}(A_{\mathbb{C}})=\bigcup_{n}\mathcal{U}_{n}(A_{\mathbb{C}})\;.$
Recall that a unitary $u\in M_{n}(A_{\mathbb{C}})$ is identified here with $\left(\begin{smallmatrix}u&0\\ 0&1\end{smallmatrix}\right)$ in $M_{n+1}(A_{\mathbb{C}})$. The identity is the class of the unit $[1]$ and the group inverse of an element $[u]$ in $K_{1}^{u}(A_{\mathbb{C}})$ is given by $u^{*}]$, since the unitary $\left(\begin{smallmatrix}u&0\\ 0&u^{*}\end{smallmatrix}\right)$ can be connected by a path of unitaries to $\left(\begin{smallmatrix}1&0\\ 0&1\end{smallmatrix}\right)$.
@@
Let $u$ be a unitary in $M_{n}(A_{\mathbb{C}})$. Let $x_{t}$ be a path of unitaries in $\mathcal{U}_{2n}({\mathbb{C}})$ from $\left(\begin{smallmatrix}0&1_{n}\\ 1_{n}&0\end{smallmatrix}\right)$ to $1_{2n}$, and let $v_{t}$ be the path
@@
Then $p_{t}=v_{t}\left(\begin{smallmatrix}1_{n}&0\\ 0&0\end{smallmatrix}\right)v_{t}^{*}$ is a projection in $C([0,1],M_{2n}(A_{\mathbb{C}}))$ with $p_{0}=p_{1}=\left(\begin{smallmatrix}1_{n}&0\\ 0&0\end{smallmatrix}\right)$. The element
@@
We refer to the proof of Theorem 10.1.3 for the details that $\Theta$ is well-defined and gives an isomorphism. ∎
@@
Let $A$ be a real unital $C^{*}$-algebra. Then $KO_{1}(A)$ is isomorphic to the group of homotopy classes in $\mathcal{U}_{\infty}(A_{\mathbb{R}})$, which is isomorphic to the group of homotopy classes in $\mathcal{U}_{\infty}(A_{\mathbb{C}})$ of unitaries $u$ that satisfy $u^{*}=u^{\tau}$.
@@
The second statement is true simply because the condition $u^{*}=u^{\tau}$ (for an element in $M_{n}(A_{\mathbb{C}})$) is equivalent to the condition that $u$ is in $M_{n}(A)$.
For the first statement, the proof is the same as in the complex case, with one extra point. Note that $\left(\begin{smallmatrix}0&1_{n}\\ 1_{n}&0\end{smallmatrix}\right)$ is not in the same homotopy class of $1_{2n}$ in $M_{2n}({\mathbb{R}})$ when $n$ is odd. However, we solve this by moving to a larger matrix algebra. Specifically, for $n$ odd
@@
is in the same homotopy class of unitaries as $1_{2n+1}$. Using this, we can produce a path $v_{t}$ of unitaries from $1_{2n+1}$ to $\mathrm{diag}(u,u^{*},1)$ in $M_{2n+1}({\mathbb{R}})$. The rest of the proof is the same as for Proposition 25.7. ∎
@@
Since $S{\mathbb{R}}=C_{o}^{{\mathbb{R}}}({\mathbb{R}})$ is not unital, we must consider these elements in the unitization
@@
where $\pi$ is evaluation at $z=1\in S^{1}$. The unitaries in the fourth column of the table shows the generators of the various non-trivial $KO_{i}$ groups in this picture.